# Dynamics in an emergent quantum-like state space generated by a nonlinear classical network

### v3

@@ VI Conclusions
We studied a specific example of a large classical system that maps, via a graph, to a state space that we call quantum like (QL) because it mimics the kinds of states typical of quantum systems. The classical system is a network of coupled phase oscillators, of the kind extensively studied in a wide variety of contexts. Such networks of coupled phase oscillators evolve according to the nonlinear dynamics associated with synchronization of the oscillator phases. We reported numerical simulations for dynamics of this system, starting from a state preparation produces a random set of initial oscillator phases. The network topology was blueprinted by a graph that generates the QL states. We used the phases of the classical system to weight the edges of the QL state graph at various time points during the dynamics, then mapped those graphs to the QL state space. Studying the state space, we constructed the density matrix in an effective basis of two coupled QL bits and showed that the mixedness of the emergent state diminished as the underlying classical system synchronized.
@@
Similarly, decoherence of an initially synchronized system was exhibited by starting the dynamics of a weakly-coupled phase oscillator system with initial conditions where all the oscillators were in phase. During that simulation, the oscillators came out of phase, and the emergent state become concomitantly mixed. The superpositions of states in the product basis collapse, evidenced by the decay of off-diagonal elements in the density matrix. Increasing mixedness and spontaneous collapse of superpositions is caused by loss of synchronization among the many classical degrees of freedom (modeled as phase oscillators). We likened the many degrees of freedom of these variables to a kind of intrinsic ‘bath’. The QL state space therefore evolved, for the chosen parameters, like an open quantum system. Based on these results it is proposed that in the opposite limit, of a strongly phase-locked classical network—that is, where couplings between phase oscillators are very large so the classical system remains synchronized—the state space can evolve unitarily. We see evidence for this regime as the purity-preserving dynamics in the equilibrium emergent states of a synchronized network.
Key conclusions of the work are that nonlinear dynamics in the classical system controls the dynamics in the state space. The dynamics can help quantum-like states be robust to noise and disorder, or they can act like an intrinsic environment to produce decoherence in the state space. We showed numerical studies of both scenarios. Nevertheless, these dynamics convey linear maps in the state space (Prop. 1). This setup suggests a mechanism by which underlying variables act to collapse entangled states in the QL system. It is speculated that the concept might inspire a plausible mechanism that, with suitable development, might explain the quantum-to-classical transition. Further, we compared QL states to quantum states. We found that during decoherence of the QL states, the off-diagonal density matrix elements decay and that this decay can be observed in any basis we choose for measurement. More surprisingly, we showed that a no-cloning theorem (that is, a state of a QL bit cannot be copied) applies not only to the QL states, but also to the underlying classical system.

### v4

@@ VII Conclusions
We studied a specific example of a large classical system that maps, via a graph, to a state space that we call quantum like (QL) because it mimics the kinds of states typical of quantum systems. The classical system is a network of coupled phase oscillators, of the kind extensively studied in a wide variety of contexts. Such networks of coupled phase oscillators evolve according to the nonlinear dynamics associated with synchronization of the oscillator phases. We reported numerical simulations for dynamics of this system, starting from a state preparation produces a random set of initial oscillator phases. The network topology was blueprinted by a graph that generates the QL states. We used the phases of the classical system at various time points during the dynamics to define dynamics of the QL states by unitary maps. Studying the state space, we constructed the density matrix in an effective basis of a composite system comprising two QL bits.
@@
When the classical system was set up so that it spontaneously synchronizes, we showed that the mixedness of the emergent state diminishes as the underlying classical system synchronizes. Once the oscillators are synchronized, the dynamics of the QL states in this equilibrium regime are purity-preserving. We also observed parameters where the converse of synchronization can happen. That is, dephasing of the QL states of an initially synchronized system was exhibited by starting the dynamics of a weakly-coupled phase oscillator system with initial conditions where all the oscillators were in phase. During that simulation, the oscillators came out of phase, and the emergent state become concomitantly mixed. The superpositions of states in the product basis collapse, evidenced by the decay of off-diagonal elements in the density matrix. We likened the many degrees of freedom of these variables to a kind of intrinsic ‘bath’. The QL state space therefore evolved, for the chosen parameters, like an open quantum system.